### v1

# Incremental Joint Learning of Depth, Pose and Implicit Scene Representation on Monocular Camera in Large-scale Scenes

Abstract
Dense scene reconstruction for photo-realistic view synthesis has various applications, such as VR/AR, autonomous vehicles. However, most existing methods have difficulties in large-scale scenes due to three core challenges: (a) inaccurate depth input. Accurate depth input is impossible to get in real-world large-scale scenes. (b) inaccurate pose estimation. Most existing approaches rely on accurate pre-estimated camera poses. (c) insufficient scene representation capability. A single global radiance field lacks the capacity to effectively scale to large-scale scenes. To this end, we propose an incremental joint learning framework, which can achieve accurate depth, pose estimation, and large-scale scene reconstruction. A vision transformer-based network is adopted as the backbone to enhance performance in scale information estimation. For pose estimation, a feature-metric bundle adjustment (FBA) method is designed for accurate and robust camera tracking in large-scale scenes. In terms of implicit scene representation, we propose an incremental scene representation method to construct the entire large-scale scene as multiple local radiance fields to enhance the scalability of 3D scene representation. Extended experiments have been conducted to demonstrate the effectiveness and accuracy of our method in depth estimation, pose estimation, and large-scale scene reconstruction.
I Introduction
In the past two decades, the demand for dense scene reconstruction has increased across various applications, including mobile robotics [1, 2], autonomous driving [3, 4, 5, 6], smart warehouse [7] and virtual reality [8, 9, 10]. Many works focused on this field have made significant progress using Neural Radiance Fields (NeRF) [11]. The original NeRF method needs the pre-compute pose input from COLMAP or SLAM methods [12, 13, 14, 15], which makes it challenging to use in real-world deployments due to the long processing time. Structure from motion (SFM) is also not robust and accurate in hand-held scenes. Some recent works are proposed to simultaneously estimate camera poses and neural implicit representation, such as BARF [16], SC-NeRF [17], and so on [18, 19, 20]. However, existing methods come with the assumptions of a fixed-size map, which renders them useless in the real-world application.
Figure 1: Qualitative results on the Tanks and Template datasets [21] on Ballroom sequence. Our method renders better novel view synthesis than existing methods.
Recent works, such as NoPe-NeRF [22] and LocalRF [23], have been proposed to improve the scene representation, which is closely related to our work. They design a joint learning approach for pose and scene representation. NoPe-NeRF also estimates the scale correction parameters during training. LocalRF decomposes the scene into several local radiance fields and progressively estimates the poses of input frames. However, both of them have difficulty in dealing with long trajectories and accumulating pose drifts. They are also sensitive to pixel-wise noise variation in longer trajectory sequences.
This paper focuses on large-scale 3D reconstruction, highlighting three key challenges: (a) inaccurate depth scale estimation, crucial for accurate geometry in neural scene representation, particularly challenging in large-scale scenes; (b) inaccurate pose estimation, which current methods struggle with due to error accumulation in large-scale scenes; (c) limited scene representation capability, where existing approaches using a single global model hinder scalability to larger scenes and longer sequences.
To this end, we propose an incremental joint learning method for depth, pose estimation, and scene implicit representation. The depth estimation network adopts a vision transformer-based network to encode input images into high-level features. The transformer backbone utilizes a global receptive field to process features at every stage, which provides a more precise estimation. For the pose estimation network, a feature-metric bundle adjustment (FBA) method is proposed for robust pose estimation. Compared to the existing BA methods, our FBA method effectively avoids the sensitivity of pixel-wise noise variation, enabling robust localization in large-scale scenes. For scene implicit representation, existing methods usually use an MLP to encode the scene, leading to poor scene representation capability. We incorporate the feature grid-based NeRF with an incremental scene representation method. We dynamically initialize local radiance fields when the camera moves to the bound of the local scene representation. The incremental scene representation method can significantly enhance the scene representation capability in large-scale scenes.
The main contributions of this paper are as follows:
  • A novel incremental joint learning method is proposed for accurate depth estimation, pose estimation, and large-scale scene reconstruction.
  • We design an FBA method for the pose estimation module. Our method extracts multi-level features with pixel-wise confidences and incorporates a coarse-to-fine BA strategy to achieve robust and accurate pose estimation in large-scale scenes.
  • An incremental scene representation method is designed, which dynamically initiates local radiance fields trained with frames within a local window. This enables accurate scene reconstruction in arbitrarily long video and large-scale scenes.
  • Showcasing superior performance on both public datasets and proprietary elderly care sample tasks datasets, ensuring real-world effectiveness. The design and code will be made open-source to benefit the communities.
II Related work
Depth estimation, pose estimation, and scene representation based on monocular images have been a major problem in robotics and computer vision. Here, we summarize some relevant works.
II-A Joint Learning of Depth and Pose
Some early works [24] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images. [25] use synthetic images to help the disparity learning. Recently, some works have focused on self-supervised learning [26, 27, 28, 29] to jointly estimate pose and depth.  [30] and [31] jointly optimize pose and depth via photometric reprojection error. [32] focus on the scale ambiguity for monocular depth estimation with normalization. However, convolutional neural networks do not have a global receptive field at every stage and fail to capture global consistency.
II-B Implicit Scene Representation
With the proposal of NeRF [11], many researchers explore combining this implicit method into various applications,such as autonomous driving [33] and object detection [34, 35, 36]. However, the original NeRF method needs groundtruth pose information as their input, which is difficult to obtain in real-world environments. They require pre-computed camera parameters obtained from SFM and SLAM algorithms [37, 38, 39]. Recently, many researchers have focused on implicit scene representation without pose prior. iNeRF [40] is the first work focused on pose estimation and scene representation. They estimate the camera pose for novel view images with a pre-trained NeRF model. BARF [16] proposes a BA optimization for camera poses with and NeRF. Some other methods [41, 42, 43, 44] combine semantic feature with NeRF to improve the scene representation and camera tracking. NoPe-NeRF [22] and LocalRF [23] are the most relevant work to ours. They jointly estimate depth, pose, and scene representation. However, they often fail in large-scale scene representation, and their pose estimation module could be more robust.
III Joint learning of depth, pose, and implicit scene representation
In this work, we propose a joint learning framework for depth, pose estimation, and implicit scene representation. The architecture of our framework is shown in Fig. 2. The system only takes in RGB images. There are three networks in our system: depth network, pose network, and scene representation network. We will introduce our three networks separately.
Figure 2: The pipeline of our system. There are three modules in our system: depth estimation module, FBA pose estimation module, and incremental scene implicit representation module. All three modules are jointly optimized during the system operation.
III-A Depth Estimation
In our depth network, we choose the vision transformer network as our backbone. Vision transformer network [45] has been proven to have strong feature learning capabilities. Most large vision models, such as Segment Anything Model [46], choose it as their backbone. Inspired by DPT [47], we leverage vision transformers in place of convolutional neural networks for dense depth prediction. We maintain an encoder-decoder structure for our depth network. For the input images $I\in H\times W\times 3$, these images are decomposed into a sequence of 2D patches, where $(H,W)$ is the height and width of the original image, $(P,P)$ is the resolution of each image patches. We use ResNet50 with group norm and weight standardization to encode these image patches. Our vision transformer module transforms these patches using sequential blocks of multi-headed self-attention (MSA).
$\hat{Z}^{l}=MSA\left(LN\left(Z^{l-1}\right)\right)+Z^{l-1}$ (1)
$Z^{l}=MLP\left(LN\left(\hat{Z}^{l}\right)\right)+\hat{Z}^{l}$
$\hat{Z}^{l+1}=MSA\left(LN\left(Z^{l}\right)\right)+Z^{l}$
$Z^{l+1}=MLP\left(LN\left(\hat{Z}^{l+1}\right)\right)+\hat{Z}^{l+1},$
where $Z^{l+1}$ is the output feature in stage $l$. LN represents layer norm operation. Then, we adopt a depth estimation head with a convolution layer and linear layer to decode the feature. The final linear layer projects this representation to a non-negative scalar that represents the inverse depth prediction for every pixel. Bilinear interpolation is used to upsample the representation. This depth network can provide dense and accurate depth predictions in large-scale scenes.
Figure 3: Qualitative results on Tanks and Temples datasets [21]. We present the scene reconstruction results of NoPe-NeRF[22], LocalRF [23], and our method with PSNR value. Our method performs better than other methods with higher-quality view synthesis results.
TABLE I: Novel view synthesis results on Tanks and Temples [21]. We use PSNR, SSIM, and LPIPS as our metrics.
Scenes BARF [16] NoPe-NeRF [22] LocalRF [23] Ours
[…]
Tanks and Temples [21] Ballroom 20.66 0.50 0.60 24.97 0.71 0.39 25.29 0.89 0.08 26.50 0.85 0.08
Barn 25.28 0.64 0.48 25.77 0.68 0.44 27.89 0.88 0.11 28.59 0.91 0.11
Church 23.17 0.62 0.52 23.60 0.67 0.46 28.07 0.91 0.07 28.82 0.92 0.08
Family 23.04 0.61 0.56 23.77 0.68 0.48 28.60 0.92 0.007 30.48 0.94 0.05
Francis 25.85 0.69 0.57 29.48 0.80 0.38 31.71 0.93 0.11 33.38 0.93 0.10
Horse 24.09 0.72 0.41 27.30 0.83 0.27 29.98 0.94 0.07 30.07 0.93 0.07
Ignatius 21.78 0.47 0.60 23.77 0.61 0.47 26.54 0.86 0.09 27.33 0.87 0.09
Museum 23.58 0.61 0.55 26.39 0.76 0.36 26.35 0.86 0.16 29.43 0.92 0.12
Mean 23.42 0.61 0.54 26.34 0.74 0.39 27.87 0.91 0.09 29.04 0.90 0.09
III-B Feature-metric BA Pose Estimation
In terms of pose estimation, we propose a FBA network. We show the architecture of our network in Fig. 2. Inspired by [48], each image is encoded into high-level features and uses FBA to get an accurate and robust pose estimation. For every consecutive frames $I_{a},I_{b}$, we use a Unet-like encoder-decoder network to extract $D_{l}$ dimensional feature map $\mathbf{F}^{l}\in\mathbb{R}^{W_{l}\times H_{l}\times D_{l}}$ at each level $l\in\{L,\ldots,1\}$. Those multi-channel image features with decreasing resolution can encode richer semantic information and larger spatial information of the image. Our network can also simultaneously estimate the confidence of each pixel to select the robust features for pose estimation.
Compared with the original BA method, this learned representation is robust to large illumination or viewpoint changes, which can provide accurate pose estimation in indoor and outdoor environments. Despite poor initial pose estimates, it can also provide meaningful gradients for successful estimation. Those joint learning methods with classical direct alignment [22, 23] operate on the original RGB image, which is not robust to large-scale scenes and long-term changes, and resorts to Gaussian image pyramids, which still largely limits the convergence to frame-to-frame tracking.
Feature Residual: We define the direct feature residual between the two consecutive frames. We optimize the relative pose$\{R,t\}$ through minimizing the feature residual. For a given feature level $l$ and each 3D point $i$ observed in two consecutive frames, we define a residual:
$\mathbf{R}_{a}^{i}=\mathbf{F}_{b}^{l}\left[\Pi\left(\mathbf{R}\mathbf{P}_{i}+\mathbf{t}\right)_{b}\right]-\mathbf{F}_{a}^{l}\left[\mathbf{p}_{a}^{i}\right]$ (2)
where $\Pi\left(\mathbf{R}\mathbf{P}_{i}+\mathbf{t}\right)_{b}$ is the projection of frame b given its current pose estimate. [·] is a lookup with sub-pixel interpolation. $F_{a}^{l},F_{b}^{l}$ are the corresponding feature maps of images a and b. The total error over N observations is:
$E_{l}(\mathbf{R},\mathbf{t})=\sum_{i,a}w_{a}^{i}\rho\left(\left\|\mathbf{R}_{a}^{i}\right\|_{2}^{2}\right)$ (3)
where $\rho(\cdot)$ is a robust cost function with derivative $\rho^{\prime}$, and $\omega_{a}^{i}$ is a per-residual weight. We use the Levenberg-Marquardt (LM) algorithm to iteratively optimize this nonlinear least-squares cost from an initial estimate. Follow [48], each pose update $\delta\in R$ is parametrized on the SE(3) manifold using its Lie algebra.
$\boldsymbol{\delta}=-(\mathbf{H}+\lambda\operatorname{diag}(\mathbf{H}))^{-1}\mathbf{J}^{\top}\mathbf{Wr}$ (4)
where $\lambda$ is the damping factor and H is the Hessian matrices. Then, the final pose can be defined by left-multiplication on the manifold.
Feature-metric BA loss: Our network is trained by minimizing the FBA of the corresponding pixel from coarse level to fine level:
$\mathcal{L}_{FBA}=\frac{1}{L}\sum_{l}\sum_{i}\left\|\Pi\left(\mathbf{R}_{l}\mathbf{P}_{i}+\mathbf{t}_{l}\right)-\Pi\left(\overline{\mathbf{R}}\mathbf{P}_{i}+\overline{\mathbf{t}}\right)\right\|_{\gamma}$ (5)
where $\gamma$ is the Huber cost. With coarse-to-fine pose estimation, low-resolution feature maps are responsible for the robustness of the pose estimation, while finer features enhance its accuracy. This feature-metric BA loss weights the supervision of the rotation and translation adaptively for each training example. It greatly enhances the accuracy and robustness of our pose network in large-scale scenes.
Figure 4: Qualitative results on Static Hikes Indoor datasets [23] with 40m trajectory and span over 1000 images. We present the scene reconstruction results of NoPe-NeRF[22], LocalRF [23], and our method in different time steps (90, 220, 390). Our incremental scene representation method significantly improves the scene reconstruction performance in large-scale scenes.
III-C Incremental Implicit Scene Representation
Existing scene implicit scene representation methods such as [17, 49, 16] can only achieve satisfactory results in small-scale scenes. They are frequently constrained by their limitations in scene representation capabilities when dealing with longer sequences and tend to converge towards local minima during the optimization process. Inspired by [23], we propose an incremental implicit scene representation. The architecture of our network is shown in Fig. 2. At the beginning, we will initialize a local neural radiance field and optimize it using the initial few frames as input. As the robot moves forward, we progressively append new frames within the local neural radiance field and iteratively refine the scene representation, pose estimation, and depth estimation. This procedure persists until we reach the boundary of the local neural radiance field. Then, we dynamically allocate new local radiance fields throughout the optimization with frames within a sliding window. This procedure will continue until all the images are registered into the network. So, the entire scene can be represented as multiple local NeRF:
$\{I_{i},D_{i}\}_{i=1}^{M}\mapsto\{\mathrm{LR}^{1}_{\theta_{1}},\mathrm{LR}^{2}_{\theta_{2}},\dots,\mathrm{LR}^{n}_{\theta_{n}}\}\mapsto\{\mathbf{c},\mathbf{\sigma}\}$ (6)
where $D_{i}$ is the estimated depth image from our depth network, $\mathbf{c}$ denotes the image color, $\sigma$ represents the volume density. Whenever the estimated camera pose exceeds the boundaries of the current local radiance fields, we cease optimizing previous poses and freeze the network parameters. This allows us to decrease memory usage by discarding supervisory frames that are no longer needed. During the optimization, we jointly optimize the depth, pose, and scene representation in the local radiance fields.
TABLE II: Pose estimation results on Tanks and Temples datasets [21]. We use ATE and RPE as our metrics.
Scenes BARF [16] NoPe-NeRF [22] LocalRF [23] Ours
[…]
Tanks and Temples [21] Ballroom 0.019 0.228 0.343 0.003 0.026 0.054 0.014 0.077 0.052 0.014 0.022 0.051
Barn 0.075 0.326 1.402 0.023 0.034 0.127 0.019 0.088 0.048 0.019 0.015 0.044
Church 0.059 0.063 0.458 0.026 0.013 0.088 0.017 0.077 0.042 0.016 0.023 0.111
Family 0.116 0.595 0.577 0.006 0.045 0.049 0.004 0.074 0.043 0.004 0.021 0.035
Francis 0.095 0.749 0.924 0.005 0.029 0.069 0.004 0.035 0.082 0.003 0.025 0.068
Horse 0.016 0.399 0.239 0.005 0.032 0.192 0.037 0.287 0.561 0.014 0.027 0.035
Ignatius 0.057 0.288 1.187 0.003 0.010 0.039 0.003 0.076 0.028 0.002 0.017 0.061
Museum 0.257 1.128 2.589 0.035 0.247 0.335 0.025 0.303 0.179 0.025 0.219 0.217
Mean 0.087 0.472 0.965 0.006 0.056 0.080 0.015 0.101 0.082 0.012 0.046 0.077
Differentiable Rendering For every input image $I$, we use the camera intrinsic matrix and camera pose to generate rays $r=x_{i}+td_{i}$. We sample points $x_{i}$ along rays. Then, we query our local radiance fields to get the color and density. $F_{\Theta}:(\mathbf{x_{i}},\mathbf{d_{i}})\rightarrow(\mathbf{c},\sigma)$. Through differentiable rendering, we can render along the rays:
$\hat{\mathbf{C}}(\mathbf{r})$ $=\sum_{i=1}^{N}T_{i}\left(1-\exp\left(-\sigma_{i}\delta_{i}\right)\right)\mathbf{c}_{i},$ (7)
$T_{i}$ $=\exp\left(-\sum_{i-1}^{j}\sigma_{j}\delta_{j}\right)$
where $\delta_{i}$ is the distance between sample points and $T_{i}$ represents the accumulated transmittance along the ray. Compared with existing scene representation methods, we have re-parameterized the scene representation for unbounded scenes similar to [50]. We use a contract function to map every point into a cubic space:
$\operatorname{contract}(\mathbf{x})=\begin{cases}\mathbf{x}&\text{ if }\|\mathbf{x}\|_{\infty}\leq 1\\ \left(2-\frac{1}{\|\mathbf{x}\|_{\infty}}\right)\left(\frac{\mathbf{x}}{\|\mathbf{x}\|_{\infty}}\right)&\text{ otherwise. }\end{cases}$ (8)
This contract function can help our method to reconstruct unbounded scenes.
III-D Incremental Joint Learning
In this section, we provide more details of the optimization of depth $D$, RGB $C$, camera poses $\{R,t\}$, and scene representation $\theta$. In order to optimize our depth network and render depth images, we propose depth loss and estimated depth function.
$\hat{\mathbf{D}}(\mathbf{r})=\sum_{i=1}^{N}T_{i}\left(1-\exp\left(-\sigma_{i}\delta_{i}\right)\right)d_{i}$ (9)
$\mathcal{L}_{d}=\left|\hat{\mathbf{D}}^{*}-\mathbf{D}^{*}\right|$ (10)
where $\hat{D}^{*}(r)$ is our render depth, $D^{*}(r)$ is our estimated depth image from our depth network. We use a pre-trained model to get depth images prior to our depth network. We define a photo-metric loss to optimize local radiance fields:
$\mathcal{L}_{p}=\|\hat{\mathbf{C}}(\mathbf{r})-\mathbf{C}(\mathbf{r})\|_{2}^{2}$ (11)
where $\hat{C}(r)$ is the render image and $C(r)$ is the input RGB image. We also add an optical flow loss between neighbor frames as they have proven to improve the accuracy and robustness in challenging scenarios.
^$\hat{\mathcal{F}}_{k\rightarrow k+1}=(u,v)-\Pi\left(\{R,t\}_{k\rightarrow k+1}\Pi^{-1}(u,v,\hat{D})\right)$ (12)
Where $\Pi^{-1}$ represents the back-projection from pixel coordinates into 3D points. $\{R,t\}_{k\rightarrow k+1}$ means the relative camera pose transformation. Then, we calculate our optical flow loss:
$\mathcal{L}_{fa}=\left\|\hat{\mathcal{F}}_{k\rightarrow k+1}-\mathcal{F}_{k\rightarrow k+1}\right\|_{1}$ (13)
This is the forward optical flow and we also calculate backward optical flow $L_{fb}$ to supervise our implicit scene representation network. The final loss of our network is :
$\mathcal{L}=\lambda_{1}\mathcal{L}_{p}+\lambda_{2}\mathcal{L}_{D}+\lambda_{3}\mathcal{L}_{FBA}+\lambda_{4}\mathcal{L}_{fa}+\lambda_{5}\mathcal{L}_{fb}$ (14)
We use the progressive optimization scheme to improve the robustness and to satisfy large-scale scenes and long sequences. We start by using a small amount of frames (for example five frames) to initiate the operation. Subsequently, we progressively introduce new frames into the optimization process in a sequential manner.
TABLE III: Novel view synthesis on Static Hikes dataset [23] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
Methods Static Hikes Indoor [23] Proprietary Dataset
[…]
BARF [16] 14.81 0.692 0.737 0.632 0.545 1.791 12.38 0.56 0.75 2.463 1.334 2.482
NoPe-NeRF [22] 16.82 0.74 0.62 0.463 0.443 1.595 20.17 0.68 0.56 0.970 0.140 1.459
LocalRF [23] 20.08 0.702 0.448 0.231 0.846 1.443 19.63 0.649 0.432 0.234 0.585 0.707
Ours 21.65 0.709 0.306 0.203 0.941 1.752 22.09 0.764 0.277 0.232 0.385 0.610
IV Experiments
In this section, we present our experimental results on different real-world datasets and conduct a comprehensive ablation study to verify the effectiveness of our method. We also gather proprietary real-world datasets using a robotic wheelchair to assess performance on mobile platforms in real-world applications.
Baselines Methods and Public Datasets. The main baseline methods we compared are BARF [16], NoPe-NeRF [22], and LocalRF [23]. All of them are joint learning methods of camera poses and implicit scene representation. We evaluate our method on the Tanks and Temples dataset [21] and Static Hikes dataset [23]. The Tanks and Temples dataset features slow-moving image sequences spanning a few meters in both indoor and outdoor scenes. We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 30 meters. We strictly follow the experimental settings of our baseline to partition the training set and the test set.
Our Dataset. An in-house proprietary dataset was collected using an intelligent wheelchair to simulate the application of NeRF in elderly care settings, as shown in Fig. 5. The dataset includes a long-distance travel scene spanning over 35 meters and consists of over 1k images. Our intelligent wheelchair features two sensors: the Realsense D435i, an RGBD camera capturing only the RGB images, and the Robosense RS-Lidar-16, aimed at obtaining groundtruth poses with a certified performance at 2 cm level accuracy [51]. We evaluate the capabilities of our depth, pose estimation, and scene representation. We use this dataset to validate the performance of our method in real-world robot operations.
Figure 5: The image above is a prototype of our robot platform, and below are the qualitative results of our method on our own real-world datasets with PSNR value. We compare the view synthesis results with NoPe-NeRF[22], LocalRF [23]. Our method performs better on real-world datasets.
Figure 6: Qualitative results of depth estimation on Tanks and Temples datasets [21]. We present the depth estimation results of NoPe-NeRF[22], LocalRF [23], and our method. The region outlined on the image is marked in red and black to signify lower accuracy, in green to signify higher accuracy and groundtruth.
Implementation Details. We present our implementation details. We select Adam optimizer[52] ($\beta=(0.9,0.999)$) for scene representation and camera tracking optimization. The color loss weighting is $\lambda_{1}=0.25$, $\lambda_{2}=0.1$ for depth and $\lambda_{3}=1.0$ for pose estimation, $\lambda_{4}=1.0$, $\lambda_{5}=1.0$ for optical flow. All training and evaluation experiments are conducted on a single NVIDIA RTX 3090 GPU.
IV-A Experimental Results
In this section, we evaluate the system in different datasets ranging from indoor scenes to large outdoor scenes on Tanks and Temples [21], Static Hikes Indoor [23] and our dataset.
Scene Reconstruction Results. We use Peak Signal-to-noise Ratio (PSNR), Structural Similarity Index Measurement (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to evaluate view synthesis results. In Table I, we present the scene reconstruction results on Tanks and Temples datasets. Our method outperforms all the baselines by a large margin. The scene reconstruction results on other datasets are shown in Table III. Our method outperforms other methods in scene representation. The incremental implicit scene representation method significantly enhances the scene representation ability and performs well on large-scale and long-sequence datasets. The qualitative results are shown in Fig. 3,4,5. In Fig. 3, We present the view synthesis results on different sequences Ignatius, Horses, Church in Tanks and Temples dataset [21]. Our method achieves higher-quality image reconstruction results. In Fig. 4, these images correspond to different timesteps (90,390,900) from up to down. We place the PSNR values of each scene in the bottom right corner. In contrast, our incremental scene representation method achieves high-quality view synthesis results compared with NoPe-NeRF and LocalRF. In Fig. 5, we show the image reconstruction results on our own dataset. Compared with other joint learning methods, our method achieves the most high-fidelity novel view results in both outdoor and indoor scenes.
Pose Estimation Results. The pose estimation results are shown in Table. II and Table. III. We use Absolute Trajectory Error (ATE), and Relative Pose Error (RPE) as the pose estimation metrics. Our pose estimation network performs better than other existing methods. The FBA method performs re-projection optimization in the high-dimensional feature level, which avoids various disturbances and noise present at the image level. The FBA network enhances the accuracy and robustness of the pose estimation in large-scale scenes. We also present the depth estimation results in Fig. 6. Compared to our method, a tree is missing from the LocalRF estimated depth, as shown in red blocks, and the depth at both bushes and trees in NoPe-NeRF is inaccurate, as shown in black blocks. Our estimated depth images are more accurate and sharper, as shown in green blocks.
Figure 7: Importance of incremental scene representation. We compare the scene reconstruction results with two types of radiance fields model (single, multiple) on Tanks and Temples dataset [21] and Static Hikes dataset [23]. In the first row, we use a single radiance field, which fails in large-scale scenes. In the second row, we use local radiance fields, which improve the representation capability and produce high-fidelity and sharper results.
TABLE IV: Ablation studies of different modules on Tanks and Temples datasets [21].
Tanks and Temples [21]
[…]
(a) w/o FBA-net 27.15 0.87 0.10 0.017 0.157 0.143
(b) w/o PO 25.87 0.84 0.15 0.016 0.152 0.141
(c) w/o $L_{d}$ 26.75 0.86 0.11 0.017 0.143 0.137
Ours 29.11 0.90 0.09 0.013 0.139 0.133
IV-B Ablation Study
In this section, we conduct sufficient ablation studies to verify the effectiveness of our designed network. We show the ablation results in Table IV. (a) is the network without a FBA network. (b) is the network without progressive optimization. (c) is the network without depth loss. For the FBA network, we remove it in Table IV. (a), and replace it with traditional pose optimization. We can see that this network improves the accuracy of the pose estimation results. For the progressive optimization network, we replace it with global optimization. We can see that the progressive network significantly improves the scene representation and pose estimation. We also present the render results between using a single global radiance field and using multiple radiance fields in Fig. 7. This indicates the importance of our incremental scene representation method. In Table (c), we remove the depth loss and the depth network is not jointly optimized with other networks. The results indicate that the depth loss and joint optimization strategy are effective for scene representation and pose estimation.
V Conclusion
In this paper, we propose a novel incremental joint learning network of depth, pose, and implicit scene representation. The proposed FBA network extracts multi-level image features and performs coarse-to-fine BA estimation for corresponding pixels and image pose, which can provide accurate pose estimation results both in indoor and outdoor scenes. An incremental scene representation method is designed for large-scale scenes and long-sequence videos. This proposed network can achieve accurate and robust scene representation for arbitrarily long-sequence scene reconstruction. Compared with other joint learning methods, our method achieves SOTA results in depth estimation, pose estimation, and implicit scene representation in large-scale indoor and outdoor scenes.

### v2

@@
 (b) inaccurate pose estimation.
-Most existing approaches rely on accurate pre-estimated camera poses.
+Existing methods are not robust enough with the growth of cumulative errors in large scenes and long sequence.
 (c) insufficient scene representation capability.
@@
 Extended experiments have been conducted to demonstrate the effectiveness and accuracy of our method in depth estimation, pose estimation, and large-scale scene reconstruction.
+The code will be open-sourced upon paper acceptance.
+Index Terms:
+Neural Radiance Fields, Incremental Joint Learning, Pose Estimation, Large-scale Scenes.
 I Introduction
-In the past two decades, the demand for dense scene reconstruction has increased across various applications, including mobile robotics [1, 2], autonomous driving [3, 4, 5, 6], smart warehouse [7] and virtual reality [8, 9, 10].
+In the past two decades, the demand for dense scene reconstruction has increased across various applications, including mobile robotics, autonomous driving.
 Many works focused on this field have made significant progress using Neural Radiance Fields (NeRF) [1].
-The original NeRF method needs the pre-compute pose input from COLMAP or SLAM methods [12, 13, 14, 15], which makes it challenging to use in real-world deployments due to the long processing time.
+The original NeRF method needs the pre-compute pose input from COLMAP or SLAM methods [2], which makes it challenging to use in real-world deployments due to the long processing time.
 Structure from motion (SFM) is also not robust and accurate in hand-held scenes.
@@
 However, existing methods come with the assumptions of a fixed-size map, which renders them useless in the real-world application.
-Figure 1: Qualitative results on the Tanks and Template datasets [21] on Ballroom sequence. Our method renders better novel view synthesis than existing methods.
+Figure 1: Qualitative results on the Tanks and Template datasets [8] on Ballroom sequence. Our method achieve better view synthesis performance than existing methods.
 Recent works, such as NoPe-NeRF [9] and LocalRF [10], have been proposed to improve the scene representation, which is closely related to our work.
@@
 NoPe-NeRF also estimates the scale correction parameters during training.
-LocalRF decomposes the scene into several local radiance fields and progressively estimates the poses of input frames.
+LocalRF progressively estimates the poses of input frames.
 However, both of them have difficulty in dealing with long trajectories and accumulating pose drifts.
-They are also sensitive to pixel-wise noise variation in longer trajectory sequences.
+Their end-to-end pose estimation are also sensitive to pixel-wise noise variation in longer trajectory sequences.
 This paper focuses on large-scale 3D reconstruction, highlighting three key challenges: (a) inaccurate depth scale estimation, crucial for accurate geometry in neural scene representation, particularly challenging in large-scale scenes; (b) inaccurate pose estimation, which current methods struggle with due to error accumulation in large-scale scenes; (c) limited scene representation capability, where existing approaches using a single global model hinder scalability to larger scenes and longer sequences.
 To this end, we propose an incremental joint learning method for depth, pose estimation, and scene implicit representation.
+These three elements are closely interconnected: depth information is crucial for geometric representation of the scene.
+It can guide the neural point sampling and enhance scene geometry consistency.
+Accurate pose estimation is also vital for scene representation.
+So, we design a joint learning framework for these three elements.
 The depth estimation network adopts a vision transformer-based network to encode input images into high-level features.
@@
 For scene implicit representation, existing methods usually use an MLP to encode the scene, leading to poor scene representation capability.
-We incorporate the feature grid-based NeRF with an incremental scene representation method.
+We propose an incremental scene representation method.
 We dynamically initialize local radiance fields when the camera moves to the bound of the local scene representation.
@@
 II-A Joint Learning of Depth and Pose
-Some early works [24] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images.
+Some early works [11, 12] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images.
 [13] use synthetic images to help the disparity learning.
-Recently, some works have focused on self-supervised learning [26, 27, 28, 29] to jointly estimate pose and depth.
-[30] and [31] jointly optimize pose and depth via photometric reprojection error.
+Recently, some works have focused on self-supervised learning to jointly estimate pose and depth.
+[14, 15, 16, 17, 18, 19] and [20] jointly optimize pose and depth via photometric reprojection error.
 [21] focus on the scale ambiguity for monocular depth estimation with normalization.
+[22] proposes a masked GAN network for unsupervised monocular depth and ego-motion estimation.
+[23] jointly perform depth and camera pose estimation and semantic scene segmentation.
 However, convolutional neural networks do not have a global receptive field at every stage and fail to capture global consistency.
+[24] incorporate channelwise attention mechanism spatialwise attention mechanism for accurate depth estimation.
 II-B Implicit Scene Representation
-With the proposal of NeRF [11], many researchers explore combining this implicit method into various applications,such as autonomous driving [33] and object detection [34, 35, 36].
-However, the original NeRF method needs groundtruth pose information as their input, which is difficult to obtain in real-world environments.
+Early Novel View Synthesis methods applied interpolations between pixels [25], later works use meshes, Multi-Plane Images, and light field [26].
+With the proposal of NeRF [1], many researchers explore combining this implicit method into various applications, such as autonomous driving [27] and robot navigation.
+[28] integrates neural radiance fields with navigation networks end-to-end as a memory structure for the first time, highlighting the potential applications of neural radiance fields in downstream control domain tasks.
+[29] design a waypoint-based task decomposition strategy which effectively enhances the reliability of learning-based navigation.
+The key idea of NeRF is to model the radiance field of a scene with a multi-layer perceptron (MLP).
+However, the representation of volume densities can not commit the geometric consistency, leading to poor surface prediction for reconstruction tasks.
+UNISURF [30] and NeuS [31] are proposed to combine world-centric 3D geometry representation with neural radiance fields.
+They replace the volume density with Signed Distance Field (SDF) values, which can further improve the surface reconstruction.
+Other methods [32, 33] use various scene geometry representation methods, such as truncated signed distance function, voxel grid.
+Figure 2: The pipeline of our system.
+There are three modules in our system: depth estimation module, Feature-metric Bundle Adjustment pose estimation module, and incremental scene implicit representation module.
+All three modules are jointly optimized during the system operation.
+II-C NeRF with Pose Optimization
+The original NeRF method needs groundtruth pose information as their input, which is difficult to obtain in real-world environments.
 They require pre-computed camera parameters obtained from SFM and SLAM algorithms [34, 35, 36].
@@
 They estimate the camera pose for novel view images with a pre-trained NeRF model.
-BARF [16] proposes a BA optimization for camera poses with and NeRF.
+NeRFmm [38] jointly optimised as learnable parameters with NeRF training, through a photometric reconstruction.
+BARF [3] proposes a coarse-to-fine Bundle Adjustment method to estimate camera poses with neural 3D representation.
 Some other methods [39, 40, 41, 42] combine semantic feature with NeRF to improve the scene representation and camera tracking.
 NoPe-NeRF [9] and LocalRF [10] are the most relevant work to ours.
-They jointly estimate depth, pose, and scene representation.
-However, they often fail in large-scale scene representation, and their pose estimation module could be more robust.
+NoPe-NeRF learns scale and shift parameters during training and incorporates undistorted monocular depth prior.
+However, they often fail in large-scale scene representation due to the limited scene representation capacity.
+Their pose estimation module are not robust and accurate in large-scale scenes, resulting in increased pose drifts.
 III Joint learning of depth, pose, and implicit scene representation
@@
 We will introduce our three networks separately.
-Figure 2: The pipeline of our system.
-There are three modules in our system: depth estimation module, FBA pose estimation module, and incremental scene implicit representation module.
-All three modules are jointly optimized during the system operation.
 III-A Depth Estimation
@@
 This depth network can provide dense and accurate depth predictions in large-scale scenes.
-Figure 3: Qualitative results on Tanks and Temples datasets [21].
-We present the scene reconstruction results of NoPe-NeRF[22], LocalRF [23], and our method with PSNR value.
-Our method performs better than other methods with higher-quality view synthesis results.
-TABLE I: Novel view synthesis results on Tanks and Temples [21]. We use PSNR, SSIM, and LPIPS as our metrics.
-Scenes BARF [16] NoPe-NeRF [22] LocalRF [23] Ours
-[…]
-Tanks and Temples [21] Ballroom 20.66 0.50 0.60 24.97 0.71 0.39 25.29 0.89 0.08 26.50 0.85 0.08
-Barn 25.28 0.64 0.48 25.77 0.68 0.44 27.89 0.88 0.11 28.59 0.91 0.11
-Church 23.17 0.62 0.52 23.60 0.67 0.46 28.07 0.91 0.07 28.82 0.92 0.08
-Family 23.04 0.61 0.56 23.77 0.68 0.48 28.60 0.92 0.007 30.48 0.94 0.05
-Francis 25.85 0.69 0.57 29.48 0.80 0.38 31.71 0.93 0.11 33.38 0.93 0.10
-Horse 24.09 0.72 0.41 27.30 0.83 0.27 29.98 0.94 0.07 30.07 0.93 0.07
-Ignatius 21.78 0.47 0.60 23.77 0.61 0.47 26.54 0.86 0.09 27.33 0.87 0.09
-Museum 23.58 0.61 0.55 26.39 0.76 0.36 26.35 0.86 0.16 29.43 0.92 0.12
-Mean 23.42 0.61 0.54 26.34 0.74 0.39 27.87 0.91 0.09 29.04 0.90 0.09
 III-B Feature-metric BA Pose Estimation
@@
 Feature Residual: We define the direct feature residual between the two consecutive frames.
-We optimize the relative pose$\{R,t\}$ through minimizing the feature residual.
+We optimize the relative pose$\{\mathbf{R},\mathbf{t}\}$ through minimizing the feature residual.
 For a given feature level $l$ and each 3D point $i$ observed in two consecutive frames, we define a residual:
 $\mathbf{r}_{a}^{i}=\mathbf{F}_{b}^{l}\left[\Pi\left(\mathbf{R}\mathbf{P}_{i}+\mathbf{t}\right)_{b}\right]-\mathbf{F}_{a}^{l}\left[\mathbf{p}_{a}^{i}\right]$ (2)
-where $\Pi\left(\mathbf{R}\mathbf{P}_{i}+\mathbf{t}\right)_{b}$ is the projection of frame b given its current pose estimate.
-[·] is a lookup with sub-pixel interpolation.
+where $\mathbf{p}_{a}^{i}=\Pi\left(\mathbf{R}\mathbf{P}_{i}+\mathbf{t}\right)_{b}$ is the projection of frame b given its current pose estimate.
+[]delimited-[][$\cdot$][ ⋅ ] is a lookup with sub-pixel interpolation.
 $F_{a}^{l},F_{b}^{l}$ are the corresponding feature maps of images a and b.
@@
 where $\lambda$ is the damping factor and H is the Hessian matrices. Then, the final pose can be defined by left-multiplication on the manifold.
+Algorithm 1 Incremental Implicit Scene Representation.
+1:  $j\leftarrow 1$; /$\ast$/ ∗ The first scene representation index/$\ast$/∗ /
+2:  $p\leftarrow 1$; /$\ast$/ ∗ First camera pose index/$\ast$/∗ /
+3:  $q\leftarrow 1$; /$\ast$/ ∗Start with the first frames $\ast$/∗ /
+4:  $\{R,T\}_{p..q}\leftarrow 1$; /$\ast$/ ∗ Initialize poses as identity/$\ast$/∗ /
+5:  $\theta_{j}\leftarrow InitializeLR()$; /$\ast$/ ∗ Initialize the first local radiance fields/$\ast$/∗ /
+6:  while $q<P$ do
+7:     while $m=0;m<100$ do
+8:        Optimize $(\theta_{j})$; /$\ast$/ ∗ Optimize the first local radiance field with the first frame/$\ast$/∗ /
+9:        $m\leftarrow m+1$;
+10:     end
+11:     while $\{R,t\}_{j}<Bound_{j}$ do
+12:        $q\leftarrow q+1$;
+13:        $\{R,T\}_{q}\leftarrow\{R,T\}_{q-1}$; /$\ast$/ ∗ Append a pose at the end of the trajectory $\ast$/∗ /
+[…]
+15:     end
+16:     if $q<P$ then
+17:        $j\leftarrow j+1$; /$\ast$/ ∗ A new local radiance field index/$\ast$/∗ /
+18:        $\theta_{j}\leftarrow InitializeLR()$ /$\ast$/ ∗ Initialize a new local radiance field/$\ast$/∗ /
+19:        $t_{j}\leftarrow t_{q}$; /$\ast$/ ∗ Centered around the last pose/$\ast$/∗ /
+20:        $p\leftarrow p+n$; /$\ast$/ ∗ Stop considering the frames in last local radiance field/$\ast$/∗ /
+21:     end if
+22:     Repeat until all frames are registered.
+23:  end
 Feature-metric BA loss: Our network is trained by minimizing the FBA of the corresponding pixel from coarse level to fine level:
@@
 It greatly enhances the accuracy and robustness of our pose network in large-scale scenes.
-Figure 4: Qualitative results on Static Hikes Indoor datasets [23] with 40m trajectory and span over 1000 images.
-We present the scene reconstruction results of NoPe-NeRF[22], LocalRF [23], and our method in different time steps (90, 220, 390).
-Our incremental scene representation method significantly improves the scene reconstruction performance in large-scale scenes.
+Figure 3: Qualitative results on Tanks and Temples datasets [8].
+We present the scene reconstruction results of NoPe-NeRF[9], LocalRF [10], and our method with PSNR value.
+Our method performs better than other methods with higher-quality view synthesis results.
+TABLE I: Novel view synthesis results on Tanks and Temples [8]. We use PSNR, SSIM, and LPIPS as our metrics.
+Scenes BARF [3] NoPe-NeRF [9] LocalRF [10] Ours
+[…]
+Tanks and Temples [8] Ballroom 20.66 0.50 0.30 24.97 0.71 0.20 25.29 0.89 0.08 26.50 0.85 0.08
+Barn 25.28 0.64 0.24 25.77 0.68 0.22 27.89 0.88 0.11 28.59 0.91 0.11
+Church 23.17 0.62 0.27 23.60 0.67 0.23 28.07 0.91 0.07 28.82 0.92 0.08
+Family 23.04 0.61 0.28 23.77 0.68 0.22 28.60 0.92 0.007 30.48 0.94 0.05
+Francis 25.85 0.69 0.28 29.48 0.80 0.18 31.71 0.93 0.11 33.38 0.93 0.10
+Horse 24.09 0.72 0.21 27.30 0.83 0.13 29.98 0.94 0.07 30.07 0.93 0.07
+Ignatius 21.78 0.47 0.30 23.77 0.61 0.23 26.54 0.86 0.09 27.33 0.87 0.09
+Museum 23.58 0.61 0.28 26.39 0.76 0.17 26.35 0.86 0.16 29.43 0.92 0.12
+Mean 23.42 0.61 0.27 26.34 0.74 0.19 27.87 0.91 0.09 29.04 0.90 0.09
 III-C Incremental Implicit Scene Representation
-Existing scene implicit scene representation methods such as [17, 49, 16] can only achieve satisfactory results in small-scale scenes.
+Existing scene implicit scene representation methods such as [4, 3] can only achieve satisfactory results in small-scale scenes.
 They are frequently constrained by their limitations in scene representation capabilities when dealing with longer sequences and tend to converge towards local minima during the optimization process.
@@
 The architecture of our network is shown in Fig.
-2.
+2 and Alg.
+1.
 At the beginning, we will initialize a local neural radiance field and optimize it using the initial few frames as input.
@@
 During the optimization, we jointly optimize the depth, pose, and scene representation in the local radiance fields.
-TABLE II: Pose estimation results on Tanks and Temples datasets [21]. We use ATE and RPE as our metrics.
-Scenes BARF [16] NoPe-NeRF [22] LocalRF [23] Ours
-[…]
-Tanks and Temples [21] Ballroom 0.019 0.228 0.343 0.003 0.026 0.054 0.014 0.077 0.052 0.014 0.022 0.051
-Barn 0.075 0.326 1.402 0.023 0.034 0.127 0.019 0.088 0.048 0.019 0.015 0.044
-Church 0.059 0.063 0.458 0.026 0.013 0.088 0.017 0.077 0.042 0.016 0.023 0.111
-Family 0.116 0.595 0.577 0.006 0.045 0.049 0.004 0.074 0.043 0.004 0.021 0.035
-Francis 0.095 0.749 0.924 0.005 0.029 0.069 0.004 0.035 0.082 0.003 0.025 0.068
-Horse 0.016 0.399 0.239 0.005 0.032 0.192 0.037 0.287 0.561 0.014 0.027 0.035
-Ignatius 0.057 0.288 1.187 0.003 0.010 0.039 0.003 0.076 0.028 0.002 0.017 0.061
-Museum 0.257 1.128 2.589 0.035 0.247 0.335 0.025 0.303 0.179 0.025 0.219 0.217
-Mean 0.087 0.472 0.965 0.006 0.056 0.080 0.015 0.101 0.082 0.012 0.046 0.077
 Differentiable Rendering For every input image $I$, we use the camera intrinsic matrix and camera pose to generate rays $r=x_{i}+td_{i}$.
@@
 This contract function can help our method to reconstruct unbounded scenes.
+Depth Guided Sampling  The estimated depth images provide valuable geometry information which can guide neural point sampling along a ray.
+We get $N_{strat}$ points for stratified sampling between the near and far planes.
+Then, $N_{surface}$ points are drawn from the Gaussian distribution determined by the depth prior.
+When the depth is not known or invalid, we use the rendered depth.
+Compared to the original methods, our approach allows for more efficient point sampling and enhances the scene representation capability of the network.
+Figure 4: Qualitative results on Static Hikes Indoor datasets [10] with long trajectory and span over 1000 images.
+We present the scene reconstruction results of NoPe-NeRF[9], LocalRF [10], and our method in different time steps (90, 220, 390, 700).
+Our incremental scene representation method significantly improves the scene reconstruction performance in large-scale scenes.
 III-D Incremental Joint Learning
@@
 $\mathcal{L}_{d}=\left|\hat{\mathbf{D}}^{*}-\mathbf{D}^{*}\right|$ (10)
-where $\hat{D}^{*}(r)$ is our render depth, $D^{*}(r)$ is our estimated depth image from our depth network.
+where $\hat{D}^{*}(r)$ is the normalized render depth, $D^{*}(r)$ is the normalized estimated depth image from our depth network.
 We use a pre-trained model to get depth images prior to our depth network.
+Since the monocular depth is not scale- and shift-invariant, we normalize the depth and estimate the scale and shift of depth:
+$t(D)=\frac{1}{N}\sum_{i=1}^{N}(D)$ (11)
+$s(D)=\frac{1}{N}\sum_{i=1}^{N}|D-t(D)|$ (12)
+We sample M pixels for the frame. Then, we normalize the depth:
+$D^{*}=\frac{D-t(D)}{s(D)}$ (13)
 We define a photo-metric loss to optimize local radiance fields:
@@
 ^$\hat{\mathcal{F}}_{k\rightarrow k+1}=(u,v)-\Pi\left(\{R,t\}_{k\rightarrow k+1}\Pi^{-1}(u,v,\hat{D})\right)$ (15)
+TABLE II: Pose estimation results on Tanks and Temples datasets [8]. We use ATE and RPE as our metrics.
+Scenes BARF [3] NoPe-NeRF [9] LocalRF [10] Ours
+[…]
+Tanks and Temples [8] Ballroom 0.019 0.228 0.343 0.003 0.026 0.054 0.014 0.077 0.052 0.014 0.022 0.051
+Barn 0.075 0.326 1.402 0.023 0.034 0.127 0.019 0.088 0.048 0.019 0.015 0.044
+Church 0.059 0.063 0.458 0.026 0.013 0.088 0.017 0.077 0.042 0.016 0.023 0.111
+Family 0.116 0.595 0.577 0.006 0.045 0.049 0.004 0.074 0.043 0.004 0.021 0.035
+Francis 0.095 0.749 0.924 0.005 0.029 0.069 0.004 0.035 0.082 0.003 0.025 0.068
+Horse 0.016 0.399 0.239 0.005 0.032 0.192 0.037 0.287 0.561 0.014 0.027 0.035
+Ignatius 0.057 0.288 1.187 0.003 0.010 0.039 0.003 0.076 0.028 0.002 0.017 0.061
+Museum 0.257 1.128 2.589 0.035 0.247 0.335 0.025 0.303 0.179 0.025 0.219 0.217
+Mean 0.087 0.472 0.965 0.006 0.056 0.080 0.015 0.101 0.082 0.012 0.046 0.077
 Where $\Pi^{-1}$ represents the back-projection from pixel coordinates into 3D points.
@@
 $\mathcal{L}_{fa}=\left\|\hat{\mathcal{F}}_{k\rightarrow k+1}-\mathcal{F}_{k\rightarrow k+1}\right\|_{1}$ (16)
-This is the forward optical flow and we also calculate backward optical flow $L_{fb}$ to supervise our implicit scene representation network. The final loss of our network is :
+This is the forward optical flow, and we also calculate backward optical flow $L_{fb}$ to supervise our implicit scene representation network. The final loss of our network is :
 $\mathcal{L}=\lambda_{1}\mathcal{L}_{p}+\lambda_{2}\mathcal{L}_{D}+\lambda_{3}\mathcal{L}_{FBA}+\lambda_{4}\mathcal{L}_{fa}+\lambda_{5}\mathcal{L}_{fb}$ (17)
@@
 Subsequently, we progressively introduce new frames into the optimization process in a sequential manner.
-TABLE III: Novel view synthesis on Static Hikes dataset [23] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
-Methods Static Hikes Indoor [23] Proprietary Dataset
+TABLE III: The view synthesis results on Tanks and temples dataset. We report the average PSNR, SSIM, LPIPS results.
 […]
-BARF [16] 14.81 0.692 0.737 0.632 0.545 1.791 12.38 0.56 0.75 2.463 1.334 2.482
-NoPe-NeRF [22] 16.82 0.74 0.62 0.463 0.443 1.595 20.17 0.68 0.56 0.970 0.140 1.459
-LocalRF [23] 20.08 0.702 0.448 0.231 0.846 1.443 19.63 0.649 0.432 0.234 0.585 0.707
-Ours 21.65 0.709 0.306 0.203 0.941 1.752 22.09 0.764 0.277 0.232 0.385 0.610
+COLMAP NeRFfacto[48] 24.17 0.65 0.194
+Mip-NeRF360[47] 27.74 0.77 0.152
+NoPe-NeRF[9] 26.43 0.74 0.184
+LocalRF[10] 28.04 0.88 0.096
+Ours 28.17 0.88 0.095
+Joint learning SCNeRF [4] 20.82 0.58 0.227
+NeRFmm [38] 22.53 0.64 0.297
+BARF [3] 23.42 0.64 0.275
+NoPe-NeRF [9] 26.34 0.74 0.194
+LocalRF [10] 27.87 0.91 0.094
+Ours 29.04 0.90 0.092
 IV Experiments
@@
 All of them are joint learning methods of camera poses and implicit scene representation.
+We also compare our method against SCNeRF [4], NeRFmm [38], NeRFfacto [48].
 We evaluate our method on the Tanks and Temples dataset [8] and Static Hikes dataset [10].
 The Tanks and Temples dataset features slow-moving image sequences spanning a few meters in both indoor and outdoor scenes.
-We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 30 meters.
+We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 100 meters.
 We strictly follow the experimental settings of our baseline to partition the training set and the test set.
@@
 Our method performs better on real-world datasets.
-Figure 6: Qualitative results of depth estimation on Tanks and Temples datasets [21].
-We present the depth estimation results of NoPe-NeRF[22], LocalRF [23], and our method.
-The region outlined on the image is marked in red and black to signify lower accuracy, in green to signify higher accuracy and groundtruth.
 Implementation Details.
@@
 All training and evaluation experiments are conducted on a single NVIDIA RTX 3090 GPU.
+Our initial learning rate are $5\times 10^{-3}$ for rotations and $5\times 10^{-4}$ for translations.
+The resolution of our local radiance fields is $64^{3}$.
+When the camera moves to the bound, we refine the local radiance fields and the camera pose for 600 iterations per frame and upsamle the resolution to $640^{3}$.
+After that, we allocate new local radiance fields and repeat this process until all frames are registered.
+TABLE IV: Novel view synthesis on Static Hikes dataset [10] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
+Methods Static Hikes Indoor [10] Our Dataset
+[…]
+BARF [3] 14.81 0.692 0.737 0.632 0.545 1.791 12.38 0.56 0.75 2.463 1.334 2.482
+NoPe-NeRF [9] 16.82 0.74 0.62 0.463 0.443 1.595 20.17 0.68 0.56 0.970 0.140 1.459
+LocalRF [10] 20.08 0.702 0.448 0.231 0.846 1.443 19.63 0.649 0.432 0.234 0.585 0.707
+Ours 21.65 0.709 0.306 0.203 0.941 1.752 22.09 0.764 0.277 0.232 0.385 0.610
 IV-A Experimental Results
@@
 Our method outperforms all the baselines by a large margin.
-The scene reconstruction results on other datasets are shown in Table III.
+In Tab.
+III, we compare our method with other view synthesis methods with COLMAP pose, which shows the proves the advantages of our joint learning system in large-scale scenes.
+The scene reconstruction results on other datasets are shown in Table IV.
 Our method outperforms other methods in scene representation.
@@
 Compared with other joint learning methods, our method achieves the most high-fidelity novel view results in both outdoor and indoor scenes.
+Figure 6: Qualitative results of depth estimation on Tanks and Temples datasets [8].
+We present the depth estimation results of NoPe-NeRF[9], LocalRF [10], and our method.
+The region outlined on the image is marked in red and black to signify lower accuracy, in green to signify higher accuracy and groundtruth.
 Pose Estimation Results.
@@
 II and Table.
-III.
+IV.
 We use Absolute Trajectory Error (ATE), and Relative Pose Error (RPE) as the pose estimation metrics.
@@
 In the second row, we use local radiance fields, which improve the representation capability and produce high-fidelity and sharper results.
-TABLE IV: Ablation studies of different modules on Tanks and Temples datasets [21].
-Tanks and Temples [21]
-[…]
-(a) w/o FBA-net 27.15 0.87 0.10 0.017 0.157 0.143
-(b) w/o PO 25.87 0.84 0.15 0.016 0.152 0.141
-(c) w/o $L_{d}$ 26.75 0.86 0.11 0.017 0.143 0.137
-Ours 29.11 0.90 0.09 0.013 0.139 0.133
+TABLE V: Ablation studies of different modules on Tanks and Temples datasets [8].
+Tanks and Temples
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ATE $\downarrow$ RPEr $\downarrow$ RPEt $\downarrow$
+(a) w/o Dep. 26.75 0.86 0.11 0.016 0.063 0.098
+(b) w/o ISR 26.57 0.84 0.15 0.015 0.065 0.101
+(c) w/o FBA 27.15 0.87 0.10 0.016 0.068 0.103
+(d) w COLMAP 28.17 0.88 0.10 0.014 0.051 0.081
+Ours 29.04 0.90 0.09 0.012 0.046 0.077
 IV-B Ablation Study
 In this section, we conduct sufficient ablation studies to verify the effectiveness of our designed network.
-We show the ablation results in Table IV.
-(a) is the network without a FBA network.
-(b) is the network without progressive optimization.
-(c) is the network without depth loss.
-For the FBA network, we remove it in Table IV.
-(a), and replace it with traditional pose optimization.
-We can see that this network improves the accuracy of the pose estimation results.
-For the progressive optimization network, we replace it with global optimization.
-We can see that the progressive network significantly improves the scene representation and pose estimation.
+We show the ablation results in Table V.
+(a) we remove the depth estimation and depth prior.
+(b) is the network without incremental scene representation.
+(c) is the network without feature-metric bundle adjustment.
+(d) is our method using pose from COLMAP.
+In Table (a), we remove the depth loss and the depth network is not jointly optimized with other networks.
+The results indicate that the depth loss and joint optimization strategy are effective for scene representation and pose estimation.
+(b) For the incremental scene representation, we replace it with global optimization.
+We can see that the incremental scene representation significantly improves the scene representation and pose estimation.
 We also present the render results between using a single global radiance field and using multiple radiance fields in Fig.
@@
 This indicates the importance of our incremental scene representation method.
-In Table (c), we remove the depth loss and the depth network is not jointly optimized with other networks.
-The results indicate that the depth loss and joint optimization strategy are effective for scene representation and pose estimation.
+(c) We remove the feature-metric bundle adjustment and replace it with traditional pose optimization.
+We can see that this network improves the accuracy of the pose estimation results.
+(d) We use the camera pose from COLMAP and fix the camera pose throughout the entire optimization.
+This demonstrates that our joint optimization method outperforms the original method based on COLMAP.
 V Conclusion

### v3

@@
 Abstract
-Dense scene reconstruction for photo-realistic view synthesis has various applications, such as VR/AR, autonomous vehicles.
-However, most existing methods have difficulties in large-scale scenes due to three core challenges: (a) inaccurate depth input.
+Dense scene reconstruction for photo-realistic view synthesis has various applications, such as VR/AR, and robotics navigation.
+Existing dense reconstruction methods are primarily designed for small room scenarios, but in practice, the scenes encountered by robots are typically large-scale environments.
+Most existing methods have difficulties in large-scale scenes due to three core challenges: (a) inaccurate depth input.
+Depth information is crucial for both scene geometry reconstruction and pose estimation.
 Accurate depth input is impossible to get in real-world large-scale scenes.
 (b) inaccurate pose estimation.
-Existing methods are not robust enough with the growth of cumulative errors in large scenes and long sequence.
+Existing methods are not robust enough with the growth of cumulative errors in large scenes and long sequences.
 (c) insufficient scene representation capability.
-A single global radiance field lacks the capacity to effectively scale to large-scale scenes.
-To this end, we propose an incremental joint learning framework, which can achieve accurate depth, pose estimation, and large-scale scene reconstruction.
-A vision transformer-based network is adopted as the backbone to enhance performance in scale information estimation.
-For pose estimation, a feature-metric bundle adjustment (FBA) method is designed for accurate and robust camera tracking in large-scale scenes.
+A single global radiance field lacks the capacity to scale effectively to large-scale scenes.
+To this end, we propose an incremental joint learning framework, which can achieve accurate depth, pose estimation, and large-scale dense scene reconstruction.
+For depth estimation, a vision transformer-based network is adopted as the backbone to enhance performance in scale information estimation.
+For pose estimation, a feature-metric bundle adjustment (FBA) method is designed for accurate and robust camera tracking in large-scale scenes and eliminates pose drift.
 In terms of implicit scene representation, we propose an incremental scene representation method to construct the entire large-scale scene as multiple local radiance fields to enhance the scalability of 3D scene representation.
-Extended experiments have been conducted to demonstrate the effectiveness and accuracy of our method in depth estimation, pose estimation, and large-scale scene reconstruction.
-The code will be open-sourced upon paper acceptance.
+In local radiance fields, we propose a tri-plane based scene representation method to further improve the accuracy and efficiency of scene reconstruction.
+We conduct extensive experiments on various datasets, including our own collected data, to demonstrate the effectiveness and accuracy of our method in depth estimation, pose estimation, and large-scale scene reconstruction.
+The code has been open-sourced on https://github.com/dtc111111/incre-dpsr.
+Note to Practitioners—In practical robot deployment scenarios, dense scene reconstruction can play a crucial role, such as in precise localization and obstacle avoidance.
+However, existing scene reconstruction methods are mainly suitable for small-room environments and struggle with larger-scale settings like long corridors or extended sequences.
+Our method introduces an incremental joint learning framework that simultaneously performs depth estimation, pose estimation, and scene reconstruction.
+Experimental results demonstrate the accuracy and effectiveness of this approach across different scene types.
 Index Terms:
-Neural Radiance Fields, Incremental Joint Learning, Pose Estimation, Large-scale Scenes.
+3D Scene Reconstruction, Incremental Joint Learning, Neural Radiance Fields, Pose Estimation, Large-scale Scenes.
+Figure 1: Qualitative results on the Tanks and Template datasets on Ballroom sequence.
+We present the visualization results of different methods’ reconstruction results in the four corners of the image.
+Our method achieve better view synthesis performance than existing methods.
 I Introduction
-In the past two decades, the demand for dense scene reconstruction has increased across various applications, including mobile robotics, autonomous driving.
+In the past two decades, the demand for scene reconstruction has increased across various applications, including mobile robotics, autonomous driving.
 Many works focused on this field have made significant progress using Neural Radiance Fields (NeRF) [1].
-The original NeRF method needs the pre-compute pose input from COLMAP or SLAM methods [2], which makes it challenging to use in real-world deployments due to the long processing time.
-Structure from motion (SFM) is also not robust and accurate in hand-held scenes.
+However, the original NeRF method is only applicable to small room scenes and cannot perform accurate reconstruction in large-scale scenes, which limits the applicability of these methods in real-world robotic scenarios.
+In addition, the original NeRF method needs the accurate pose input, which makes it challenging to use in real-world deployments due to the inability to obtain accurate poses in practical robot operating environments.
+Pre-processing methods, such as Structure from motion (SFM), is also not robust and accurate in large-scale scenes.
 Some recent works are proposed to simultaneously estimate camera poses and neural implicit representation, such as BARF [2], SC-NeRF [3], and so on [4, 5, 6].
-However, existing methods come with the assumptions of a fixed-size map, which renders them useless in the real-world application.
-Figure 1: Qualitative results on the Tanks and Template datasets [8] on Ballroom sequence. Our method achieve better view synthesis performance than existing methods.
-Recent works, such as NoPe-NeRF [9] and LocalRF [10], have been proposed to improve the scene representation, which is closely related to our work.
-They design a joint learning approach for pose and scene representation.
-NoPe-NeRF also estimates the scale correction parameters during training.
-LocalRF progressively estimates the poses of input frames.
+Recent works, such as NoPe-NeRF [7] and LocalRF [8], have been proposed to further improve the scene representation and pose estimation.
+NoPe-NeRF estimates the scale correction parameters during training to improve scene representation.
+LocalRF progressively estimates the poses of input frames to improve the scalability.
 However, both of them have difficulty in dealing with long trajectories and accumulating pose drifts.
-Their end-to-end pose estimation are also sensitive to pixel-wise noise variation in longer trajectory sequences.
-This paper focuses on large-scale 3D reconstruction, highlighting three key challenges: (a) inaccurate depth scale estimation, crucial for accurate geometry in neural scene representation, particularly challenging in large-scale scenes; (b) inaccurate pose estimation, which current methods struggle with due to error accumulation in large-scale scenes; (c) limited scene representation capability, where existing approaches using a single global model hinder scalability to larger scenes and longer sequences.
-To this end, we propose an incremental joint learning method for depth, pose estimation, and scene implicit representation.
+Their RGB loss-based pose estimation are also sensitive to pixel-wise noise variation in long trajectory sequences.
+In this paper, we focus on large-scale 3D reconstruction without pose information, utilizing only monocular camera input.
+We highlight three key challenges in large-scale scenes: (a) inaccurate depth scale estimation Depth information is crucial for accurate geometry in neural scene representation, particularly in large-scale scenes.
+However, existing methods neglect the importance of depth information in large-scale scene reconstruction and also fail to produce accurate depth.
+(b) inaccurate pose estimation Current methods struggle with pose estimation due to the error accumulation in large-scale scenes; (c) limited scene representation capability Existing approaches using a single, global model, which hinders scalability to larger scenes and longer sequences.
+To this end, we propose an incremental joint learning framework for depth, pose estimation, and scene implicit representation in large-scale scenes.
 These three elements are closely interconnected: depth information is crucial for geometric representation of the scene.
@@
 We dynamically initialize local radiance fields when the camera moves to the bound of the local scene representation.
+In local radiance fields, unlike the coordinate-based MLPs representation, we incorporate triplane with three corresponding axes XYZ.
+Our feature plane-based scene representation method effectively improves both the accuracy and efficiency of scene representation.
+Compared to the feature grid approach, our method substantially reduces the space complexity from cubic ($\mathcal{O}(n^{3})$) to square ($\mathcal{O}(n)$).
 The incremental scene representation method can significantly enhance the scene representation capability in large-scale scenes.
 The main contributions of this paper are as follows:
-  • A novel incremental joint learning method is proposed for accurate depth estimation, pose estimation, and large-scale scene reconstruction.
-  • We design an FBA method for the pose estimation module.
+  • A novel incremental joint learning framework is proposed for accurate depth estimation, pose estimation, and large-scale scene reconstruction.
+  • We design an feature-metric bundle adjustment method for the pose estimation module.
 Our method extracts multi-level features with pixel-wise confidences and incorporates a coarse-to-fine BA strategy to achieve robust and accurate pose estimation in large-scale scenes.
   • An incremental scene representation method is designed, which dynamically initiates local radiance fields trained with frames within a local window.
+We also incorporate triplane based scene representation with three corresponding axes for accurate and efficient scene reconstruction.
 This enables accurate scene reconstruction in arbitrarily long video and large-scale scenes.
@@
 II Related work
-Depth estimation, pose estimation, and scene representation based on monocular images have been a major problem in robotics and computer vision. Here, we summarize some relevant works.
+Depth estimation [9], pose estimation [10, 11, 12], and scene representation based on monocular images have been a major problem in robotic perception and navigation [13].
+Here, we summarize some relevant works.
 II-A Joint Learning of Depth and Pose
-Some early works [11, 12] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images.
+Depth and pose are closely interconnected components.
+Some early works [14] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images.
 [15] use synthetic images to help the disparity learning.
 Recently, some works have focused on self-supervised learning to jointly estimate pose and depth.
-[14, 15, 16, 17, 18, 19] and [20] jointly optimize pose and depth via photometric reprojection error.
+[16, 17, 18, 19] and [20] jointly optimize pose and depth via photometric reprojection error.
 [21] focus on the scale ambiguity for monocular depth estimation with normalization.
-[22] proposes a masked GAN network for unsupervised monocular depth and ego-motion estimation.
-[23] jointly perform depth and camera pose estimation and semantic scene segmentation.
 However, convolutional neural networks do not have a global receptive field at every stage and fail to capture global consistency.
-[24] incorporate channelwise attention mechanism spatialwise attention mechanism for accurate depth estimation.
 II-B Implicit Scene Representation
-Early Novel View Synthesis methods applied interpolations between pixels [25], later works use meshes, Multi-Plane Images, and light field [26].
-With the proposal of NeRF [1], many researchers explore combining this implicit method into various applications, such as autonomous driving [27] and robot navigation.
+With the proposal of NeRF [1], many researchers explore combining this implicit scene representation into various applications, such as autonomous driving [22, 23, 24, 25], and robot navigation.
 [26] integrates neural radiance fields with navigation networks end-to-end as a memory structure for the first time, highlighting the potential applications of neural radiance fields in downstream control domain tasks.
@@
 They replace the volume density with Signed Distance Field (SDF) values, which can further improve the surface reconstruction.
-Other methods [32, 33] use various scene geometry representation methods, such as truncated signed distance function, voxel grid.
 Figure 2: The pipeline of our system.
@@
 The original NeRF method needs groundtruth pose information as their input, which is difficult to obtain in real-world environments.
-They require pre-computed camera parameters obtained from SFM and SLAM algorithms [34, 35, 36].
+They require pre-computed camera parameters obtained from SFM and SLAM algorithms [30, 31, 32, 33, 34, 35].
 Recently, many researchers have focused on implicit scene representation without pose prior.
@@
 BARF [2] proposes a coarse-to-fine Bundle Adjustment method to estimate camera poses with neural 3D representation.
-Some other methods [39, 40, 41, 42] combine semantic feature with NeRF to improve the scene representation and camera tracking.
+[38, 39, 40, 41, 42] combine semantic segementation methods [43, 44] with NeRF to improve the scene representation and camera tracking.
+Some other methods [45, 46, 47, 48] use multiple submaps to represent the scene.
 NoPe-NeRF [7] and LocalRF [8] are the most relevant work to ours.
 NoPe-NeRF learns scale and shift parameters during training and incorporates undistorted monocular depth prior.
-However, they often fail in large-scale scene representation due to the limited scene representation capacity.
+However, they often fail in large-scale scene representation due to the inaccurate depth estimation, the accumulated pose drift, and the limited scene representation capacity.
 Their pose estimation module are not robust and accurate in large-scale scenes, resulting in increased pose drifts.
@@
 2.
-The system only takes in RGB images.
-There are three networks in our system: depth network, pose network, and scene representation network.
+The system only takes in monocular RGB images $I\in H\times W\times 3$.
+There are three networks in our system: depth estimation network, pose estimation network, and scene representation network.
 We will introduce our three networks separately.
@@
 In our depth network, we choose the vision transformer network as our backbone.
-Vision transformer network [43] has been proven to have strong feature learning capabilities.
+Vision transformer network [49] has shown strong feature learning capabilities.
 Most large vision models, such as Segment Anything Model [50], choose it as their backbone.
-Inspired by DPT [45], we leverage vision transformers in place of convolutional neural networks for dense depth prediction.
+Inspired by this, we leverage vision transformers in place of convolutional neural networks for dense depth prediction.
 We maintain an encoder-decoder structure for our depth network.
@@
 III-B Feature-metric BA Pose Estimation
-In terms of pose estimation, we propose a FBA network.
+In terms of pose estimation, we propose a Feature-metric bundle adjustment network.
 We show the architecture of our network in Fig.
@@
 We use the Levenberg-Marquardt (LM) algorithm to iteratively optimize this nonlinear least-squares cost from an initial estimate.
-Follow [46], each pose update $\delta\in R$ is parametrized on the SE(3) manifold using its Lie algebra.
+Follow [51], each pose update is parametrized on the SE(3) manifold using its Lie algebra.
 $\boldsymbol{\delta}=-(\mathbf{H}+\lambda\operatorname{diag}(\mathbf{H}))^{-1}\mathbf{J}^{\top}\mathbf{Wr}$ (4)
-where $\lambda$ is the damping factor and H is the Hessian matrices. Then, the final pose can be defined by left-multiplication on the manifold.
+where $\lambda$ is the damping factor and H is the Hessian matrices. Then, the final pose can be defined by left-multiplication on the manifold SE(3).
 Algorithm 1 Incremental Implicit Scene Representation.
 1:  $j\leftarrow 1$; /$\ast$/ ∗ The first scene representation index/$\ast$/∗ /
-2:  $p\leftarrow 1$; /$\ast$/ ∗ First camera pose index/$\ast$/∗ /
+2:  $p\leftarrow 1$; /$\ast$/ ∗ First camera pose index, P denotes the length of sequence/$\ast$/∗ /
 3:  $q\leftarrow 1$; /$\ast$/ ∗Start with the first frames $\ast$/∗ /
@@
 6:  while $q<P$ do
-7:     while $m=0;m<100$ do
-8:        Optimize $(\theta_{j})$; /$\ast$/ ∗ Optimize the first local radiance field with the first frame/$\ast$/∗ /
-9:        $m\leftarrow m+1$;
-10:     end
-11:     while $\{R,t\}_{j}<Bound_{j}$ do
-12:        $q\leftarrow q+1$;
-13:        $\{R,T\}_{q}\leftarrow\{R,T\}_{q-1}$; /$\ast$/ ∗ Append a pose at the end of the trajectory $\ast$/∗ /
+7:     $m\leftarrow 0$
+8:     while $m<100$ do
+9:        Optimize $(\theta_{j})$; /$\ast$/ ∗ Optimize the first local radiance field with the first frame/$\ast$/∗ /
+10:        $m\leftarrow m+1$;
+11:     end
+12:     while $\{R,t\}_{j}<Bound_{j}$ do
+13:        $q\leftarrow q+1$;
+14:        $\{R,T\}_{q}\leftarrow\{R,T\}_{q-1}$; /$\ast$/ ∗ Append a pair of pose R,T at the end of the trajectory $\ast$/∗ /
 […]
-15:     end
-16:     if $q<P$ then
-17:        $j\leftarrow j+1$; /$\ast$/ ∗ A new local radiance field index/$\ast$/∗ /
-18:        $\theta_{j}\leftarrow InitializeLR()$ /$\ast$/ ∗ Initialize a new local radiance field/$\ast$/∗ /
-19:        $t_{j}\leftarrow t_{q}$; /$\ast$/ ∗ Centered around the last pose/$\ast$/∗ /
-20:        $p\leftarrow p+n$; /$\ast$/ ∗ Stop considering the frames in last local radiance field/$\ast$/∗ /
-21:     end if
-22:     Repeat until all frames are registered.
-23:  end
+16:     end
+17:     if $q<P$ then
+18:        $j\leftarrow j+1$; /$\ast$/ ∗ A new local radiance field index/$\ast$/∗ /
+19:        $\theta_{j}\leftarrow InitializeLR()$ /$\ast$/ ∗ Initialize a new local radiance field/$\ast$/∗ /
+20:        $t_{j}\leftarrow t_{q}$; /$\ast$/ ∗ Centered around the last pose/$\ast$/∗ /
+21:        $p\leftarrow p+n$; /$\ast$/ ∗ Stop considering the frames in last local radiance field/$\ast$/∗ /
+22:     end if
+23:     Repeat until all frames are registered.
+24:  end
 Feature-metric BA loss: Our network is trained by minimizing the FBA of the corresponding pixel from coarse level to fine level:
 $\mathcal{L}_{FBA}=\frac{1}{L}\sum_{l}\sum_{i}\left\|\Pi\left(\mathbf{R}_{l}\mathbf{P}_{i}+\mathbf{t}_{l}\right)-\Pi\left(\overline{\mathbf{R}}\mathbf{P}_{i}+\overline{\mathbf{t}}\right)\right\|_{\gamma}$ (5)
-where $\gamma$ is the Huber cost.
+where $\gamma$ is the Huber cost, L denotes the different feature level.
 With coarse-to-fine pose estimation, low-resolution feature maps are responsible for the robustness of the pose estimation, while finer features enhance its accuracy.
@@
 It greatly enhances the accuracy and robustness of our pose network in large-scale scenes.
+Figure 3: We illustrate the differences between our sampling method and original sampling method. Our method makes better use of the estimated depth information to guide the ray sampling process.
 Figure 4: Qualitative results on Tanks and Temples datasets [52].
@@
 Our method performs better than other methods with higher-quality view synthesis results.
-TABLE I: Novel view synthesis results on Tanks and Temples [8]. We use PSNR, SSIM, and LPIPS as our metrics.
-Scenes BARF [3] NoPe-NeRF [9] LocalRF [10] Ours
-[…]
-Tanks and Temples [8] Ballroom 20.66 0.50 0.30 24.97 0.71 0.20 25.29 0.89 0.08 26.50 0.85 0.08
-Barn 25.28 0.64 0.24 25.77 0.68 0.22 27.89 0.88 0.11 28.59 0.91 0.11
-Church 23.17 0.62 0.27 23.60 0.67 0.23 28.07 0.91 0.07 28.82 0.92 0.08
-Family 23.04 0.61 0.28 23.77 0.68 0.22 28.60 0.92 0.007 30.48 0.94 0.05
-Francis 25.85 0.69 0.28 29.48 0.80 0.18 31.71 0.93 0.11 33.38 0.93 0.10
-Horse 24.09 0.72 0.21 27.30 0.83 0.13 29.98 0.94 0.07 30.07 0.93 0.07
-Ignatius 21.78 0.47 0.30 23.77 0.61 0.23 26.54 0.86 0.09 27.33 0.87 0.09
-Museum 23.58 0.61 0.28 26.39 0.76 0.17 26.35 0.86 0.16 29.43 0.92 0.12
-Mean 23.42 0.61 0.27 26.34 0.74 0.19 27.87 0.91 0.09 29.04 0.90 0.09
 III-C Incremental Implicit Scene Representation
@@
 During the optimization, we jointly optimize the depth, pose, and scene representation in the local radiance fields.
+TABLE I: Novel view synthesis results on Tanks and Temples [52]. We use PSNR, SSIM, and LPIPS as our metrics.
+scenes BARF[2] NoPe-NeRF [7] LocalRF [8] Flow-NeRF [53] RA-NeRF [54] Ours
+PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
+Tanks and Temples Ballroom 20.66 0.50 0.30 24.97 0.71 0.20 25.29 0.89 0.08 28.83 0.86 0.24 30.25 N/A 0.11 26.50 0.85 0.08
+Barn 25.28 0.64 0.24 25.77 0.68 0.22 27.89 0.88 0.11 28.53 0.78 0.35 25.99 N/A 0.38 28.59 0.91 0.11
+Church 23.17 0.62 0.27 23.60 0.67 0.23 28.07 0.91 0.07 28.27 0.83 0.28 32.05 N/A 0.08 28.82 0.92 0.08
+Family 23.04 0.61 0.28 23.77 0.68 0.22 28.60 0.92 0.007 29.40 0.85 0.29 30.46 N/A 0.15 30.48 0.94 0.05
+Francis 25.85 0.69 0.28 29.48 0.80 0.18 31.71 0.93 0.11 30.63 0.83 0.33 32.04 N/A 0.24 33.38 0.93 0.10
+Horse 24.09 0.72 0.21 27.30 0.83 0.13 29.98 0.94 0.07 28.57 0.86 0.23 28.26 N/A 0.18 30.07 0.93 0.07
+Ignatius 21.78 0.47 0.30 23.77 0.61 0.23 26.54 0.86 0.09 26.25 0.73 0.35 27.77 N/A 0.25 27.83 0.87 0.09
+Museum 23.58 0.61 0.28 26.39 0.76 0.17 26.35 0.86 0.16 29.43 0.85 0.27 31.59 N/A 0.10 29.43 0.92 0.12
+mean 23.42 0.61 0.27 26.34 0.74 0.19 27.87 0.91 0.09 28.73 0.82 0.29 29.80 N/A 0.18 29.84 0.90 0.09
+Local Scene Representation Voxel grid-based architectures [55] are the mainstream in NeRF-based systems.
+However, they struggle with cubical memory growing and space complexity.
+In our local scene representation, we factorize the local radiance fields to compact components and reduce the space complexity from cubic ($\mathcal{O}(n^{3})$) to square ($\mathcal{O}(n)$).
+The 3D feature grid $\mathcal{G}$ can be factorized as with Vector-Matrix (VM) decomposition:
+$\mathcal{G}=\sum_{r=1}^{R}\mathbf{v}_{r}^{X}\circ\mathbf{M}_{r}^{YZ}+\mathbf{v}_{r}^{Y}\circ\mathbf{M}_{r}^{XZ}+\mathbf{v}_{r}^{Z}\circ\mathbf{M}_{r}^{XY}$ (7)
+where $\mathbf{v}_{r}^{X}$,$\mathbf{v}_{r}^{Y}$, $\mathbf{v}_{r}^{Z}$ are the vectors corresponding to XYZ axis.
+Our factorization-based model efficiently computes the feature vector for each voxel at a low cost, requiring only a single value per XYZ-mode vector or matrix factor.
+Additionally, we incorporate a tri-linear interpolation method to achieve a continuous fields.
+In our local neural radiance field, we can just interpolate the vector/matirx factor for the corresponding modes.
+Instead of recovering eight individual tensor elements for trilinear interpolation, we directly compute the interpolated value, significantly reducing computation and memory costs during runtime.
 Differentiable Rendering For every input image $I$, we use the camera intrinsic matrix and camera pose to generate rays $r=x_{i}+td_{i}$.
@@
 $T_{i}$ $=\exp\left(-\sum_{i-1}^{j}\sigma_{j}\delta_{j}\right)$
-where $\delta_{i}$ is the distance between sample points and $T_{i}$ represents the accumulated transmittance along the ray.
+where $\delta_{i}$ is the distance between sample points, $T_{i}$ represents the accumulated transmittance along the ray, and $N$ denotes the number of sample points along the ray.
 Compared with existing scene representation methods, we have re-parameterized the scene representation for unbounded scenes similar to [56].
@@
 Compared to the original methods, our approach allows for more efficient point sampling and enhances the scene representation capability of the network.
+We illustrate the differences between our sampling method and original sampling method in Fig.
+3.
 Figure 5: Qualitative results on Static Hikes Indoor datasets [8] with long trajectory and span over 1000 images.
@@
 $s(D)=\frac{1}{N}\sum_{i=1}^{N}|D-t(D)|$ (13)
-We sample M pixels for the frame. Then, we normalize the depth:
+We sample N pixels for the frame. Then, we normalize the depth:
 $D^{*}=\frac{D-t(D)}{s(D)}$ (14)
@@
 We use the progressive optimization scheme to improve the robustness and to satisfy large-scale scenes and long sequences.
+This joint optimization framework simultaneously leverages 2D photometric consistency from RGB images and 3D structural information from depth, enforcing coherence between 2D appearance and 3D geometry.
+This synergy leads to enhanced accuracy in depth estimation, pose prediction, and overall scene reconstruction.
 We start by using a small amount of frames (for example five frames) to initiate the operation.
 Subsequently, we progressively introduce new frames into the optimization process in a sequential manner.
-TABLE III: The view synthesis results on Tanks and temples dataset. We report the average PSNR, SSIM, LPIPS results.
+TABLE III: The depth estimation accuracy on the Tanks and Temples dataset [52] and indoor datasets [8], using Depth L1, RMSE, Absolute Relative Error (Abs Rel), and Square Relative Error (Sq Rel) as metrics.
+Scenes NoPe-NeRF [7] LocalRF [8] Ours
+[…]
+[…]
+Depth-L1 [cm] $\downarrow$ 159.57 176.79 158.77
+RMSE [cm] $\downarrow$ 171.50 191.45 171.69
+[…]
+[…]
+Depth-L1$\downarrow$ 187.27 153.88 136.2
+[…]
+TABLE IV: The scene reconstruction results on Tanks and temples dataset. We report the average PSNR, SSIM, LPIPS results.
 […]
@@
 Baselines Methods and Public Datasets.
-The main baseline methods we compared are BARF [3], NoPe-NeRF [9], and LocalRF [10].
+The main baseline methods we compared are BARF [2], NoPe-NeRF [7], LocalRF [8], Flow-NeRF [53], RA-NeRF [54].
 All of them are joint learning methods of camera poses and implicit scene representation.
 We also compare our method against SCNeRF [3], NeRFmm [37], NeRFfacto [57].
-We evaluate our method on the Tanks and Temples dataset [8] and Static Hikes dataset [10].
+We evaluate our method on the Tanks and Temples dataset [52] (from 4x4m to 8x8m) and Static Hikes dataset [8] (15x15m).
 The Tanks and Temples dataset features slow-moving image sequences spanning a few meters in both indoor and outdoor scenes.
-We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 100 meters.
+We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 100 meters (30x30m).
 We strictly follow the experimental settings of our baseline to partition the training set and the test set.
@@
 An in-house proprietary dataset was collected using an intelligent wheelchair to simulate the application of NeRF in elderly care settings, as shown in Fig.
-5.
-The dataset includes a long-distance travel scene spanning over 35 meters and consists of over 1k images.
+6 and Fig.
+7.
+We collected datasets and conducted experiments in both laboratory environments and real hospital scenarios.
+The dataset includes a long-distance travel scene spanning over one hundred meters and consists of over 1k images.
 Our intelligent wheelchair features two sensors: the Realsense D435i, an RGBD camera capturing only the RGB images, and the Robosense RS-Lidar-16, aimed at obtaining groundtruth poses with a certified performance at 2 cm level accuracy [58].
@@
 Figure 6: The image above is a prototype of our robot platform, and below are the qualitative results of our method on our own real-world datasets with PSNR value.
-We compare the view synthesis results with NoPe-NeRF[9], LocalRF [10].
+We compare the scene reconstruction results with NoPe-NeRF[7], LocalRF [8].
 Our method performs better on real-world datasets.
+Figure 7: This image shows the reconstruction results of different methods in a real robot platform within a hospital scenario.
+We compare the scene reconstruction results with NoPe-NeRF[7], LocalRF [8].
+Our approach adapts better to data from various scenes, achieving more accurate scene reconstruction and pose estimation.
 Implementation Details.
@@
 We select Adam optimizer[59] ($\beta=(0.9,0.999)$) for scene representation and camera tracking optimization.
-The color loss weighting is $\lambda_{1}=0.25$, $\lambda_{2}=0.1$ for depth and $\lambda_{3}=1.0$ for pose estimation, $\lambda_{4}=1.0$, $\lambda_{5}=1.0$ for optical flow.
+The color loss weighting is $\lambda_{1}=0.25$, $\lambda_{2}=0.1$ for depth, $\lambda_{3}=1.0$ for pose estimation, and $\lambda_{4}=1.0$, $\lambda_{5}=1.0$ for optical flow.
 All training and evaluation experiments are conducted on a single NVIDIA RTX 3090 GPU.
@@
 After that, we allocate new local radiance fields and repeat this process until all frames are registered.
-TABLE IV: Novel view synthesis on Static Hikes dataset [10] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
+TABLE V: Novel view synthesis on Static Hikes dataset [8] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
 Methods Static Hikes Indoor [8] Our Dataset
@@
 Scene Reconstruction Results.
-We use Peak Signal-to-noise Ratio (PSNR), Structural Similarity Index Measurement (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to evaluate view synthesis results.
+We use Peak Signal-to-noise Ratio (PSNR), Structural Similarity Index Measurement (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to evaluate scene reconstruction results.
 In Table I, we present the scene reconstruction results on Tanks and Temples datasets.
-Our method outperforms all the baselines by a large margin.
+Our method outperforms all the baselines by a large margin, thanks to our incremental scene representation method and efficient tri-plane representation.
 In Tab.
-III, we compare our method with other view synthesis methods with COLMAP pose, which shows the proves the advantages of our joint learning system in large-scale scenes.
-The scene reconstruction results on other datasets are shown in Table IV.
+IV, we compare our method with other scene reconstruction methods with/without COLMAP pose, which shows the proves the advantages of our joint learning system in large-scale scenes.
+The scene reconstruction results on Static Hikes and our own datasets are shown in Table V.
 Our method outperforms other methods in scene representation.
@@
 The qualitative results are shown in Fig.
-3,4,5.
+4,5,6.
 In Fig.
-3, We present the view synthesis results on different sequences Ignatius, Horses, Church in Tanks and Temples dataset [8].
+4, we present the scene reconstruction results on different sequences Ignatius, Horses, Church in Tanks and Temples dataset [52].
 Our method achieves higher-quality image reconstruction results.
 In Fig.
-4, these images correspond to different timesteps (90, 390, 900) from up to down.
+5, these images correspond to different timesteps (90, 390, 900) from up to down.
 We place the PSNR values of each scene in the bottom right corner.
-In contrast, our incremental scene representation method achieves high-quality view synthesis results compared with NoPe-NeRF and LocalRF.
+In contrast, our incremental scene representation method achieves high-quality scene reconstruction results compared with NoPe-NeRF and LocalRF.
 In Fig.
-5, we show the image reconstruction results on our own dataset.
+6 and Fig.
+7, we show the image reconstruction results in our own dataset.
 Compared with other joint learning methods, our method achieves the most high-fidelity novel view results in both outdoor and indoor scenes.
-Figure 6: Qualitative results of depth estimation on Tanks and Temples datasets [8].
-We present the depth estimation results of NoPe-NeRF[9], LocalRF [10], and our method.
-The region outlined on the image is marked in red and black to signify lower accuracy, in green to signify higher accuracy and groundtruth.
+Figure 8: Visualization of pose estimation results on Tanks and Temples datasets [52].
+We present the pose estimation results of BARF [2], NoPe-NeRF[7], LocalRF [8], and our method.
+We visualize the Absolute Trajectory Error ATE of different methods.
+The color bar on the right represents the ATE pose error.
+Figure 9: Qualitative depth estimation results on the family sequence from the Tanks and Temples dataset [52]. We present the depth estimation results of NoPe-NeRF[7], LocalRF [8], and our method.
+Figure 10: Qualitative results of depth estimation on Indoor datasets [52]. We present the depth estimation results of NoPe-NeRF[7], LocalRF [8], and our method.
 Pose Estimation Results.
@@
 II and Table.
-IV.
+V.
 We use Absolute Trajectory Error (ATE), and Relative Pose Error (RPE) as the pose estimation metrics.
 Our pose estimation network performs better than other existing methods.
+We also visualize the ATE error of different methods in Tanks datasets.
 The FBA method performs re-projection optimization in the high-dimensional feature level, which avoids various disturbances and noise present at the image level.
 The FBA network enhances the accuracy and robustness of the pose estimation in large-scale scenes.
-We also present the depth estimation results in Fig.
-6.
-Compared to our method, a tree is missing from the LocalRF estimated depth, as shown in red blocks, and the depth at both bushes and trees in NoPe-NeRF is inaccurate, as shown in black blocks.
+Depth Estimation Results.
+We present the depth estimation results in Fig.
+9,10.
+In Fig.
+9, compared to our method, a tree is missing from the LocalRF estimated depth, as shown in red blocks, and the depth at both bushes and trees in NoPe-NeRF is inaccurate, as shown in black blocks.
 Our estimated depth images are more accurate and sharper, as shown in green blocks.
+We present quantitative results in Tab.
+III.
+Our depth estimation performance is better than other NeRF-Based methods thanks to the joint learning methods.
 Figure 11: Importance of incremental scene representation.
@@
 In the second row, we use local radiance fields, which improve the representation capability and produce high-fidelity and sharper results.
-TABLE V: Ablation studies of different modules on Tanks and Temples datasets [8].
+TABLE VI: Ablation studies of different modules on Tanks and Temples datasets [52].
 Tanks and Temples
@@
 In this section, we conduct sufficient ablation studies to verify the effectiveness of our designed network.
-We show the ablation results in Table V.
+We show the ablation results in Table VI.
 (a) we remove the depth estimation and depth prior.
@@
 We also present the render results between using a single global radiance field and using multiple radiance fields in Fig.
-7.
+11.
 This indicates the importance of our incremental scene representation method.
@@
 In this paper, we propose a novel incremental joint learning network of depth, pose, and implicit scene representation.
-The proposed FBA network extracts multi-level image features and performs coarse-to-fine BA estimation for corresponding pixels and image pose, which can provide accurate pose estimation results both in indoor and outdoor scenes.
+The proposed FBA network extracts multi-level image features and performs coarse-to-fine bundle adjustment estimation for corresponding pixels and image pose, which can provide accurate pose estimation results both in indoor and outdoor scenes.
 An incremental scene representation method is designed for large-scale scenes and long-sequence videos.
@@
 Compared with other joint learning methods, our method achieves SOTA results in depth estimation, pose estimation, and implicit scene representation in large-scale indoor and outdoor scenes.
+Future work Currently, our experiments are primarily conducted in relatively large-scale indoor environments.
+As future work, we plan to extend our approach to even larger, city-scale scenes.
+To ensure better global consistency in such expansive environments, it is also necessary to incorporate techniques such as loop closure detection into our framework.